# Homomorphism, substructure and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

### v1

@@ IV.1 Tricritical Ising to Ising
The tricritical Ising CFT has six primary fields and one can decompose it to the tensor product of Ising and Fibonacci fusion rule. Hence, we denote the model as,
@@
In both fusion rules, $I$ represents the corresponding identity object. The remarkable point of this representation is $Z_{2}$ group-like action of $\psi$ and this $Z_{2}$ structure is called simple current in physics. Corresponding to this $Z_{2}$ like property, one can introduce the $Z_{2}$ charge. In this manuscript, we study $Z_{2}$ charge preserving homomorphism or RG flow satisfying the t’Hooft anomaly matching condition. In the above fusion rule, $\{I,\psi\}$ are uncharged and $\{\sigma\}$ is charged.
@@
In this setting, one can construct ring homomorphism when the homomorphism acts on the object $\tau$ canonically. Because of the theft anomaly matching condition, the action of $\rho$ to $\tau$ should be in the following form
@@
Hence by inserting the condition coming from the t’Hooft anomaly matching condition, one can obtain the following equations,
@@
It might be surprising for the readers in theoretical physics, that only the preservation of symmetry coming from symmetry-preserving domain wall condition cannot fix the homomorphism. We further introduce the assumption on the preservation of the fermion parity structure. In this setting, one can restrict the solutions to those with $\beta=0$. We can obtain a single solution by imposing the positivity of the coefficient. The solution corresponds to the RG domain wall is,
@@
Before going to the next example, we remark that we have assumed that $\psi$ is mapped to $\psi^{\prime}$, but one can do the same analysis by replacing the chirality of $\psi$ and $\psi^{\prime}$. Corresponding to the combinations of chiral or antichiral extension, there exists a series of homomorphisms. In condensed matter theory, this appears as particle-hole conjugate symmetry[37, 38] and the above discussion provides an evident algebraic expression. Corresponding phenomena happen in general $Z_{2}$ symmetric models including other examples in this work.
@@ IV.2 $SU(2)_{1}\times SU(2)_{1}\rightarrow\text{Majorana}$
Here, we study a simple model which plays a significant role in the study of SPT in condensed matter. We strudy the ring iosomorphism between $SU(2)_{1}\times SU(2)_{1}$ to $Z_{2}$ extension of $SU(2)_{2}$. We denote the chiral objects in $SU(2)_{1}\times SU(2)_{1}$ as,
@@
Because $SU(2)_{2}$ satisfy the Ising fusion rule we denote the theory as $\{I,\psi,\sigma\}$. It has been well established that the $Z_{2}$ extension of this chiral Ising fusion rule forms double semion algebra[117].
@@
where one can represent $J^{[i]}=j^{[i]}\overline{j^{[i]}}$. To avoid complication, we drop the upper indices of the identity operators.
@@
It should be stressed that $\sigma^{\prime}_{\text{Bulk},+}\times\sigma^{\prime}_{\text{Bulk},-}$ is vanishing and one needs to introduce two sets of models when considering its lattice realization[238, 239, 240, 241]. This structure has appeared in fermionic string theory, for example in [117], but it has never captured sufficient attention in the fields. This suggests the flows of the sectors $J^{[1]}$ or $J^{[2]}$ do not behave well formally, and the $Z_{2}$ symmetric or $Z_{2}$ antisymmetric sectors $(J^{[1]}\pm J^{[2]})/\sqrt{2}$ are more useful. The corresponding structure can be seen when studying the subalgebra of $SU(2)_{1}\times SU(2)_{1}$.
@@ IV.3 $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$
Here we study the homomorphism induced from the flow $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$. We use the same notation in the previous subsection and we distinguish the simple currents of each $SU(2)_{1}$ by writing them as $J^{[i]}$, $i=1,2,3$. In $SU(2)_{3}$ there are four primary fields with spins ranging from 00 to $\frac{3}{2}$, which we list in ascending order of spin as ${I^{\prime},x^{\prime},\tau^{\prime},J\textquoteright}$ ’. Here, $\tau^{\prime}$ is the primary field that satisfies the Fibonacci fusion rule $\tau^{\prime}\times\tau^{\prime}=1+\tau^{\prime}$, and $J\textquoteright$ ’ is the simple current of $SU(2)_{3}$. Here we study the homomorphism preserving $Z_{2}$ structures $\{I,J^{[1]}J^{[2]}J^{[3]}\}$ and $\{I,J^{\prime}\}$.
@@
and regard this as a suitable linear combination. In this representation, we have assumed two conditions, preservation of $Z_{2}$ group ring structure and t’Hooft anomaly matching condition as in the previous cases. To determine the specific values of $\alpha^{ij}$ and $\beta^{ij}$, we use the fusion rules among the $J^{[i]}$. Specifically, we exploit $I=I\times I=(J^{[i]})^{2}(J^{[j]})^{2}=(J^{[i]}J^{[j]})^{2}$. Together with the previous equation and nontrivial fusion rule $\tau^{\prime}\times\tau^{\prime}=I^{\prime}+\tau^{\prime}$ results in the following form,
@@
which is consistent. We note that contrary to the naive intuition from the fact that $\tau^{\prime}$ is nonabelian anyon, above $I^{\prime}/\sqrt{5}-2\tau/\sqrt{5}$ is invertible. One can also check the invertibility of $\tau^{\prime}$ itself by calculating $\tau^{\prime}\times(\tau^{\prime}-I^{\prime})=I^{\prime}$. In summary, the admissible solutions are either $(\alpha^{ij},\beta^{ij})=(1,0)$ for every $i\neq j$, or the pattern in which two out of the three unordered pairs $(i,j)$ take $\bigl{(}\tfrac{1}{\sqrt{5}},-\tfrac{2}{\sqrt{5}}\bigr{)}$ while the remaining pair takes $(1,0)$. It is remarkable that when assuming the homomorphism is surjective, the appearance of the fractional and negative coefficients is inevitable in this case. When one prohibits the appearance of such an unusual coefficient, the homomorphism is isomorphic to the homomorphism $\{SU(2)_{1}\}^{3}\rightarrow SU(2)_{1}$ and this is consistent with the observation in [62].
@@ IV.4 $\{SU(2)_{1}\}^{4}\to SU(2)_{4}$: puzzle of charged sector
Next, we determine the homomorphism corresponding to the flow $\{SU(2)_{1}\}^{4}$to $SU(2)_{4}$ in uncharged $Z_{2}$ sector. We keep the notation for $SU(2)_{1}$ as before and label each copy by the upper indices $[i]$ and take the $Z_{2}$ simple current as $J=\prod_{i=1}^{4}J^{[i]}$. The theory $SU(2)_{4}$ has five chiral primary fields, labeled by their spin as $\phi^{\prime}_{0},\ \phi^{\prime}_{1/2},\ \phi^{\prime}_{1},\ \phi^{\prime}_{3/2},\ \phi^{\prime}_{2}$ whose fusion rules are listed in the table below. The remarkable point of these two theories is the conformal dimension $1$ of the $Z_{2}$ currents $J$ and $J^{\prime}$ and this means that the theories are classified as the simplest bosonic model in the corresponding series of models.
@@
$2$ 00
$\phi_{2}(=J\textquoteright)$ ’ ) is the $\mathbb{Z}_{2}$ simple current of the theory. By computing the $Z_{2}$ charge of each operator with respect to $J$ or $J\textquoteright$ ’, we can restrict the operators that should be connected by the homomorphism. In the present case, $\phi_{1/2}$ and $\phi_{3/2}$ belong to the $Z_{2}$ uncharged sector and the others belong to the $Z_{2}$ charged sector. One must note, as in the Majorana-fermion case, that $Z_{2}$ invariant operators split after the $Z_{2}$ simple current extension. In the present case the $\mathbb{Z}_{2}$-invariant sector is $\phi^{\prime}_{1}$, which therefore splits into $\phi^{\prime}_{1,0}$ and $\phi^{\prime}_{1,1}$ where the lower index 00, $1$ corresponds to $Z_{2}$ fermion parity. Even after the $Z_{2}$ extension, the number of operators in the uncharged and charged sectors is significantly different and this is in sharp contrast to the Majorana fermion case with the fermionic T-duality (or a variant of supersymmetry in older literature[242]). Moreover, in this case, the $Z_{2}$ invariant sector is uncharged.
@@
where $=$ is ring isomorphism and $\Phi^{\prime}$ is the objects or anyons in nonchiral fusion ring[233, 232, 155] (or spherical fusion category). For a conventional reason, we changed the lower index of $\phi^{\prime}_{1}$ to $1,0$ in $\Phi^{\prime}_{1,0}$. The $Z_{2}$ extension can be achieved by multiplying the chiral $Z_{2}$ structure to this fusion ring. For example, one can obtain $\phi^{\prime}_{0}\overline{\phi^{\prime}_{2}}=J^{\prime}\Phi^{\prime}_{0}$. Because of the extension, it is necessary to introduce the nontrivial object $\Phi^{\prime}_{1,1}=J^{\prime}\Phi^{\prime}_{1,0}$ and this corresponds to a disorder operator.
By considering the anyon condensation process, we can obtain the algebraic data of the symmetry topological field theory (SymTFT) $\{\Psi\}$ which is ring isomorphic to the extended chiral algebra. For simplicity, we concentrate on uncharged sectors which form subalgebra and denote the condensation as follows,
@@
We replace the $\Psi^{\prime}$ with the $\phi^{\prime}$ because of the ring isomorphism (or the CFT/TQFT correspondence or CFT/SymTFT correspondence). By evaluating the charge of each operator with respect to $J\textquoteright$ ’, we find that a product of two $SU(2)_{1}$ simple currents such as $J^{\prime[i]}J^{\prime[j]}$ is mapped to a linear combination of $\phi^{\prime}_{0},\ \phi^{\prime}_{1,0},\ \phi^{\prime}_{1,1},\ \phi^{\prime}_{2}$. Thus we set
@@
must hold. We can find the solutions by dividing into four cases according to whether $\beta_{0}^{ij}$ and $\beta_{1}^{ij}$ are both zero, only one is zero, or neither is zero. The answers are the solutions in which exactly one of $\alpha^{ij},\gamma^{ij}$ is ±$\pm 1$, and the ten non-trivial solutions
@@
According to this consistency condition, the solutions ±$\pm\frac{1}{3}\phi_{0}\mp\frac{2\sqrt{2}}{3}\phi_{1,0}$ and ±$\pm\frac{2}{9}\phi_{0}\pm\frac{\sqrt{2}}{9}(\phi_{1,0}+5\phi_{1,1})\mp\frac{5}{9}\phi_{2}$ are excluded. This is because multiplying this expression by any other solution cannot make the coefficient of $\phi_{0}$ vanish. Therefore, the essentially non-trivial solutions are these six in total.
@@
However, because of the relation $\rho(J^{[k]}J^{[k]})=\rho(I)=\phi^{\prime}_{0}$ and the fusion rule between $\phi^{\prime}_{1/2}$ and $\phi^{\prime}_{3/2}$, it is impossible to obtain a solution. Hence regardless of the close connection between the $\{SU(2)_{1}\}^{4}$ and $SU(2)_{4}$(see [248], for example), it is impossible to obtain a surjective ring homomorphism between them at this stage. We expect that this puzzle can be resolved by considering combinations of other nonanomalous $Z_{2}$ extensions of $\{SU(2)_{1}\}^{4}$ and the corresponding noninvertible extension of $SU(2)_{4}$ but studies on such noninvertible extension are still under development. Hence, we leave this puzzle as an open problem.
We also remark that by generalizing the arguments for the $SU(2)_{2}\rightarrow\text{Majorana}$ case, it might be possible to obtain the ring homomorphism by taking the subring of the UV theory containing charged sector. For example, the charged sector formed by $2$ dimensional basis $(\sum_{i}J^{[i]})/4$ and $(\sum_{i}JJ^{[i]})/4$ will be a reasonable charged subsector constructing the UV subalgebra.
@@
In this work, we studied the explicit construction of ring homomorphism between fusion rings which determines the possible IR symmetry systematically. Even in the well-established models connected by massless RG flows, a nontrivial structure appears. To obtain the homomorphism more generally, systematic studies on the ideal structure of the model are necessary, as we have stressed in the section. II. When the ideal is commutative, the existence of eigenvalue 00 is guaranteed and its existence in lattice models seems promissive, in principle (However, we note a more involved problem of emanant symmetry in [139, 140, 141]). In other words, if the ideal is not commutative, this implies difficulty in obtaining the IR theory by projection and the IR symmetry will become more difficult to realize in a lattice model without taking some limit. Further studies of the ideal structure will be fundamentally important for the classification and construction of general TOs. In this work, we have restricted our attention to fusion rings appearing in $2$ dimensional CFT or $2+1$ dimensional TQFT, but the same construction of RG by ideal should be applicable to higher dimensional systems. From this view, nontrivial transformations of UV and IR symmetry as in [249] are worth further studies.
@@
We also note a related research direction toward coupled models with extended symmetries. By using the folding trick[12], it is known that the construction of extended model $\mathbf{F}_{\text{UV}\times\text{IR}}$ with $\mathbf{A}_{(1)}\otimes\mathbf{A}_{(2)}\subset\mathbf{F}_{\text{UV}\times\text{IR}}$ plays a fundamental role also for the construction of homomorphism[33]. In a $Z_{N}$ symmetric case, the extension is constructed by studying a common subgroup of $\mathbf{A}_{(1)}$ and $\mathbf{A}_{(2)}$. The anomaly-free object in these theories is the integer spin simple current[73, 74, 75, 76, 77, 78, 79, 80] and its nonchiral generalizations[86, 69] and the existence of this structure enables one to obtain the extended theory $\mathbf{F}_{\text{UV}\times\text{IR}}$. Hence, one expects the same strategy for the common subalgebra structure, but the gauging or extension for the product of this common subalgebra has not been studied as far as we know. As a natural extension from $Z_{N}$ case, the nonsimple current in the coupled model forms a nonsimple current with (half) integer spin after taking the Deligne product (This condition corresponds to single-validness of the wavefunctions in TOs. The term nonsimple current apppeared in the study of heterotic string theories. This indicates nontrivial extension or insertion of duality defects as in[129, 250, 251]). For this purpose, recent progress on the orbifolding procedure for the noninvertible symmetry [103, 104] might be useful. However, we stress that it is necessary to invent the corresponding extension which is the inverse operation of the orbifolding. After the extension, the representation of the system will become intrinsically anyonic[252, 184].
@@
Finally, we make a comment on the interpretation of projection in the manuscript based on more recent literature on open quantum systems. In this manuscript, we have restricted our attention to static quantum phases and their transition, but by interpreting this projection as measurement, our construction of the projection will produce a kind of treatable measurement-induced quantum phase transition[210, 211, 212] satisfying $c$ and $g$-theorem[253, 254, 255]. Further studies on the interpretation of the ideal structure in a quantum many-body system and its realization in an experimental setting are interesting research directions. Moreover, by interpreting the ring with a negative or fractional coefficient as a category, premodular fusion categories will appear and they are expected to describe TOs in mixed states[256, 257, 258].

### v2

@@ IV.1 Tricritical Ising to Ising
The tricritical Ising CFT has six primary fields, and one can decompose it into the tensor product of Ising and Fibonacci fusion rules. Hence, we denote the model as,
@@
In both fusion rules, $I$ represents the corresponding identity object. The remarkable point of this representation is the $Z_{2}$ group-like action of $\psi$, and this $Z_{2}$ structure is called a simple current in physics. Corresponding to this $Z_{2}$ like property, one can introduce the $Z_{2}$ charge. In this manuscript, we study $Z_{2}$ charge-preserving homomorphism or RG flow satisfying the t’Hooft anomaly matching condition. In the above fusion rule, $\{I,\psi\}$ are uncharged and $\{\sigma\}$ is charged.
@@
In this setting, one can construct a ring homomorphism when the homomorphism acts on the object $\tau$ canonically. Because of the theft anomaly matching condition, the action of $\rho$ to $\tau$ should be in the following form
@@
Hence, by inserting the condition coming from the t’Hooft anomaly matching condition, one can obtain the following equations,
@@
It might be surprising for the readers in theoretical physics that only the preservation of symmetry coming from the symmetry-preserving domain wall condition cannot fix the homomorphism. We further introduce the assumption on the preservation of the fermion parity structure. In this setting, one can restrict the solutions to those with $\beta=0$. We can obtain a single solution by imposing the positivity of the coefficient. The solution corresponding to the RG domain wall is,
@@
Before going to the next example, we remark that we have assumed that $\psi$ is mapped to $\psi^{\prime}$, but one can do the same analysis by replacing the chirality of $\psi$ and $\psi^{\prime}$. Corresponding to the combinations of chiral or antichiral extension, there exists a series of homomorphisms. In condensed matter theory, this appears as particle-hole conjugate symmetry[36, 37] and the above discussion provides an evident algebraic expression. Corresponding phenomena happen in general $Z_{2}$ symmetric models, including other examples in this work.
@@ IV.2 $SU(2)_{1}\times SU(2)_{1}\rightarrow\text{Majorana}$
Here, we study a simple model that plays a significant role in the study of SPT in condensed matter. We strudy the ring iosomorphism between $SU(2)_{1}\times SU(2)_{1}$ to $Z_{2}$ extension of $SU(2)_{2}$. We denote the chiral objects in $SU(2)_{1}\times SU(2)_{1}$ as,
@@
Because $SU(2)_{2}$ satisfy the Ising fusion rule we denote the theory as $\{I,\psi,\sigma\}$. It has been well established that the $Z_{2}$ extension of this chiral Ising fusion rule forms a double semion algebra[119].
@@
where one can represent $J^{[i]}=j^{[i]}\overline{j^{[i]}}$. To avoid complications, we drop the upper indices of the identity operators.
@@
It should be stressed that $\sigma^{\prime}_{\text{Bulk},+}\times\sigma^{\prime}_{\text{Bulk},-}$ is vanishing and one needs to introduce two sets of models when considering its lattice realization[241, 242, 243, 244]. This structure has appeared in fermionic string theory, for example, in [119], but it has never captured sufficient attention in the fields. This suggests the flows of the sectors $J^{[1]}$ or $J^{[2]}$ do not behave well formally, and the $Z_{2}$ symmetric or $Z_{2}$ antisymmetric sectors $(J^{[1]}\pm J^{[2]})/\sqrt{2}$ are more useful. The corresponding structure can be seen when studying the subalgebra of $SU(2)_{1}\times SU(2)_{1}$.
@@ IV.3 $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$
Here we study the homomorphism induced from the flow $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$. We use the same notation in the previous subsection, and we distinguish the simple currents of each $SU(2)_{1}$ by writing them as $J^{[i]}$, $i=1,2,3$. In $SU(2)_{3}$ there are four primary fields with spins ranging from 0 to $\frac{3}{2}$, which we list in ascending order of spin as $I^{\prime},x^{\prime},\tau^{\prime},J\textquoteright$}. Here, $\tau^{\prime}$ is the primary field that satisfies the Fibonacci fusion rule $\tau^{\prime}\times\tau^{\prime}=1+\tau^{\prime}$, and $J\textquoteright$ is the simple current of $SU(2)_{3}$. Here we study the homomorphism preserving $Z_{2}$ structures $\{I,J^{[1]}J^{[2]}J^{[3]}\}$ and $\{I,J^{\prime}\}$.
@@
and regard this as a suitable linear combination. In this representation, we have assumed two conditions, preservation of the $Z_{2}$ group ring structure and t’Hooft anomaly matching condition as in the previous cases. To determine the specific values of $\alpha^{ij}$ and $\beta^{ij}$, we use the fusion rules among the $J^{[i]}$. Specifically, we exploit $I=I\times I=(J^{[i]})^{2}(J^{[j]})^{2}=(J^{[i]}J^{[j]})^{2}$. Together with the previous equation and nontrivial fusion rule $\tau^{\prime}\times\tau^{\prime}=I^{\prime}+\tau^{\prime}$ results in the following form,
@@
which is consistent. We note that contrary to the naive intuition from the fact that $\tau^{\prime}$ is a nonabelian anyon, above $I^{\prime}/\sqrt{5}-2\tau/\sqrt{5}$ is invertible. One can also check the invertibility of $\tau^{\prime}$ itself by calculating $\tau^{\prime}\times(\tau^{\prime}-I^{\prime})=I^{\prime}$. In summary, the admissible solutions are either $(\alpha^{ij},\beta^{ij})=(1,0)$ for every $i\neq j$, or the pattern in which two out of the three unordered pairs $(i,j)$ take $\bigl{(}\tfrac{1}{\sqrt{5}},-\tfrac{2}{\sqrt{5}}\bigr{)}$ while the remaining pair takes $(1,0)$. It is remarkable that when assuming the homomorphism is surjective, the appearance of the fractional and negative coefficients is inevitable in this case. When one prohibits the appearance of such an unusual coefficient, the homomorphism is isomorphic to the homomorphism $\{SU(2)_{1}\}^{3}\rightarrow SU(2)_{1}$ and this is consistent with the observation in [61].
@@ IV.4 $\{SU(2)_{1}\}^{4}\to SU(2)_{4}$: puzzle of charged sector
Next, we determine the homomorphism corresponding to the flow $\{SU(2)_{1}\}^{4}$ $\to$ $SU(2)_{4}$ in uncharged $Z_{2}$ sector. We keep the notation for $SU(2)_{1}$ as before and label each copy by the upper indices $[i]$ and take the $Z_{2}$ simple current as $J=\prod_{i=1}^{4}J^{[i]}$. The theory $SU(2)_{4}$ has five chiral primary fields, labeled by their spin as $\phi^{\prime}_{0},\ \phi^{\prime}_{1/2},\ \phi^{\prime}_{1},\ \phi^{\prime}_{3/2},\ \phi^{\prime}_{2}$ whose fusion rules are listed in the table below. The remarkable point of these two theories is the conformal dimension $1$ of the $Z_{2}$ currents $J$ and $J^{\prime}$, and this means that the theories are classified as the simplest bosonic model in the corresponding series of models.
@@ 11 $0+1+2$ $(1/2)+(3/2)$ 11
22 0
$\phi_{2}(=J\textquoteright)$ is the $\mathbb{Z}_{2}$ simple current of the theory. By computing the $Z_{2}$ charge of each operator concerning $J$ or $J\textquoteright$, we can restrict the operators that should be connected by the homomorphism. In the present case, $\phi_{1/2}$ and $\phi_{3/2}$ belong to the $Z_{2}$ uncharged sector and the others belong to the $Z_{2}$ charged sector. One must note, as in the Majorana-fermion case, that $Z_{2}$ invariant operators split after the $Z_{2}$ simple current extension. In the present case the $\mathbb{Z}_{2}$-invariant sector is $\phi^{\prime}_{1}$, which therefore splits into $\phi^{\prime}_{1,0}$ and $\phi^{\prime}_{1,1}$ where the lower index 0, 11 corresponds to $Z_{2}$ fermion parity. Even after the $Z_{2}$ extension, the number of operators in the uncharged and charged sectors is significantly different, and this is in sharp contrast to the Majorana fermion case with the fermionic T-duality (or a variant of supersymmetry in older literature[245]). Moreover, in this case, the $Z_{2}$ invariant sector is uncharged.
@@
where == is ring isomorphism and $\Phi^{\prime}$ is the objects or anyons in nonchiral fusion ring[236, 235, 157] (or spherical fusion category). For a conventional reason, we changed the lower index of $\phi^{\prime}_{1}$ to $1,0$ in $\Phi^{\prime}_{1,0}$. The $Z_{2}$ extension can be achieved by multiplying the chiral $Z_{2}$ structure to this fusion ring. For example, one can obtain $\phi^{\prime}_{0}\overline{\phi^{\prime}_{2}}=J^{\prime}\Phi^{\prime}_{0}$. Because of the extension, it is necessary to introduce the nontrivial object $\Phi^{\prime}_{1,1}=J^{\prime}\Phi^{\prime}_{1,0}$, and this corresponds to a disorder operator.
By considering the anyon condensation process, we can obtain the algebraic data of the symmetry topological field theory (SymTFT) $\{\Psi\}$[250], which is ring isomorphic to the extended chiral algebra. For simplicity, we concentrate on uncharged sectors which form a subalgebra and denote the condensation as follows,
@@
We replace the $\Psi^{\prime}$ with the $\phi^{\prime}$ because of the ring isomorphism (or the CFT/TQFT correspondence or CFT/SymTFT correspondence). By evaluating the charge of each operator with respect to $J\textquoteright$, we find that a product of two $SU(2)_{1}$ simple currents such as $J^{\prime[i]}J^{\prime[j]}$ is mapped to a linear combination of $\phi^{\prime}_{0},\ \phi^{\prime}_{1,0},\ \phi^{\prime}_{1,1},\ \phi^{\prime}_{2}$. Thus we set
@@
must hold. We can find the solutions by dividing into four cases according to whether $\beta_{0}^{ij}$ and $\beta_{1}^{ij}$ are both zero, only one is zero, or neither is zero. The answers are the solutions in which exactly one of $\alpha^{ij},\gamma^{ij}$ is $\pm 1$, and the ten non-trivial solutions
@@
According to this consistency condition, the solutions $\pm\frac{1}{3}\phi_{0}\mp\frac{2\sqrt{2}}{3}\phi_{1,0}$ and $\pm\frac{2}{9}\phi_{0}\pm\frac{\sqrt{2}}{9}(\phi_{1,0}+5\phi_{1,1})\mp\frac{5}{9}\phi_{2}$ are excluded. This is because multiplying this expression by any other solution cannot make the coefficient of $\phi_{0}$ vanish. Therefore, the essentially non-trivial solutions are these six in total.
@@
However, because of the relation $\rho(J^{[k]}J^{[k]})=\rho(I)=\phi^{\prime}_{0}$ and the fusion rule between $\phi^{\prime}_{1/2}$ and $\phi^{\prime}_{3/2}$, it is impossible to obtain a solution. Hence, regardless of the close connection between the $\{SU(2)_{1}\}^{4}$ and $SU(2)_{4}$(see [252], for example), it is impossible to obtain a surjective ring homomorphism between them at this stage. We expect that this puzzle can be resolved by considering combinations of other nonanomalous $Z_{2}$ extensions of $\{SU(2)_{1}\}^{4}$ and the corresponding noninvertible extension of $SU(2)_{4}$ but studies on such noninvertible extension are still under development. Hence, we leave this puzzle as an open problem.
We also remark that by generalizing the arguments for the $SU(2)_{2}\rightarrow\text{Majorana}$ case, it might be possible to obtain the ring homomorphism by taking the subring of the UV theory containing the charged sector. For example, the charged sector formed by 22 22-dimensional basis $(\sum_{i}J^{[i]})/4$ and $(\sum_{i}JJ^{[i]})/4$ will be a reasonable charged subsector constructing the UV subalgebra.
@@
In this work, we studied the explicit construction of a ring homomorphism between fusion rings that determines the possible IR symmetry systematically. Even in the well-established models connected by massless RG flows, a nontrivial structure appears. To obtain the homomorphism more generally, systematic studies on the ideal structure of the model are necessary, as we have stressed in the section. II. When the ideal is commutative, the existence of an eigenvalue 0 is guaranteed, and its existence in lattice models seems promising, in principle (However, we note a more involved problem of emanant symmetry in [141, 142, 143]). In other words, if the ideal is not commutative, this implies difficulty in obtaining the IR theory by projection, and the IR symmetry will become more difficult to realize in a lattice model without taking some limit. Further studies of the ideal structure will be fundamentally important for the classification and construction of general TOs. In this work, we have restricted our attention to fusion rings appearing in 22 22-dimensional CFT or $2+1$ dimensional TQFT, but the same construction of RG by ideal should be applicable to higher-dimensional systems. From this view, nontrivial transformations of UV and IR symmetry, as in [253, 254], are worth further study.
@@
We also note a related research direction toward coupled models with extended symmetries. By using the folding trick[11], it is known that the construction of extended model $\mathbf{F}_{\text{UV}\times\text{IR}}$ with $\mathbf{A}_{(1)}\otimes\mathbf{A}_{(2)}\subset\mathbf{F}_{\text{UV}\times\text{IR}}$ plays a fundamental role also for the construction of homomorphism[32]. In a $Z_{N}$ symmetric case, the extension is constructed by studying a common subgroup of $\mathbf{A}_{(1)}$ and $\mathbf{A}_{(2)}$. The anomaly-free object in these theories is the integer spin simple current[72, 73, 74, 75, 76, 77, 78, 79] and its nonchiral generalizations[85, 68] and the existence of this structure enables one to obtain the extended theory $\mathbf{F}_{\text{UV}\times\text{IR}}$. Hence, one expects the same strategy for the common subalgebra structure, but the gauging or extension for the product of this common subalgebra has not been studied as far as we know. As a natural extension of the $Z_{N}$ case, the nonsimple current in the coupled model forms a nonsimple current with (half) integer spin after taking the Deligne product (This condition corresponds to single-validness of the wavefunctions in TOs. The term nonsimple current appeared in the study of heterotic string theories. This indicates nontrivial extension or insertion of duality defects as in[131, 255, 256]). For this purpose, recent progress on the orbifolding procedure for the noninvertible symmetry [102, 103] might be useful. However, we stress that it is necessary to invent the corresponding extension, which is the inverse operation of the orbifolding. After the extension, the representation of the system will become intrinsically anyonic[257, 186].
@@
Finally, we make a comment on the interpretation of projection in the manuscript based on more recent literature on open quantum systems. In this manuscript, we have restricted our attention to static quantum phases and their transition, but by interpreting this projection as a measurement, our construction of the projection will produce a kind of treatable measurement-induced quantum phase transition[213, 214, 215] satisfying $c$ and $g$-theorem[258, 259, 260]. Further studies on the interpretation of the ideal structure in a quantum many-body system and its realization in an experimental setting are interesting research directions. Moreover, by interpreting the ring with a negative or fractional coefficient as a category, premodular fusion categories will appear, and they are expected to describe TOs in mixed states[261, 262, 263].